OpenStudy (anonymous):

What is the equation, in slope-intercept form, of the line that has a slope of –3 and passes through the point (0, 5)?

OpenStudy (anonymous):

put it in point-slope first: y-y1=m(x-x1) y-5=-3(x-0) y-5=-3x +5 +5 y=-3x+5

OpenStudy (anonymous):

Does this help @Dequavia123

OpenStudy (anonymous):

yess